Power dissipated across the 80 resistor in the circuit shown here is 2 watt. The power dissipated in watt units across the 30 resistor is



        (a)    1.0


        (b)    0.5


        (c)    3.0


        (d)    2.0


Grade:upto college level

1 Answers

Aditi Chauhan
10 years ago

(c)

Power = V. I = 12R

 

Potential over 80 = Ri2 = 8 ×0.5  = 4V

This is the potential over parallel branch.

So,

        Power of 3W = i1\2R = 1 × 1 × 3 = 3W